A thermometer which was not accurately calibrated, indicated 0.5% at the lower fixed point and 106 at the upper fixed point. what is the time temperature when the thermometer registers 63.4c


1
Expert's answer
2021-03-18T16:52:40-0400

The thermometer had 0.5°C at 0°C real temperature and 106°C at real 100°C. So, we can plot a graph assuming linear dependency, but it is better to find a function that connects the wrong values with real ones. Assume wrong values are along y-axis and correct are along x-axis. First, find the equation:


"\frac{y-y_1}{y_2-y_1}=\frac{x-x_1}{x_2-x_1},\\\space\\
\frac{y-0.5}{106-0.5}=\frac{x-0}{100-0},\\\space\\
x=\frac{100}{105.5}(y-0.5)."

So, 63.4°C at the wrong scale (y) correspond to correct value (x) of


"x=\frac{100}{105.5}(63.4-0.5)=59.6°\text{C}."